OpenStudy (anonymous):

(-3+-6) ________ (6-9)

OpenStudy (anonymous):

I got -6 on the calculator it says 3.

OpenStudy (anonymous):

Please help me.

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

9 divided by negative 3

OpenStudy (anonymous):

Why -3, why is my answer incorrect?

OpenStudy (anonymous):

No, it would be: \[ \frac{-3+(-6)}{6-9}=\frac{-9}{-3}=\frac{9}{3}=3 \]

OpenStudy (anonymous):

3 yeah I'm wrong hahah

OpenStudy (anonymous):

Umm, please explain.

OpenStudy (anonymous):

Adding two negatives results in the sum of the values, negated. So, we have: \[ -1+-3=-(1+3)=-(4)=-4 \]

OpenStudy (anonymous):

wait please.

OpenStudy (anonymous):

I really do not get it......

OpenStudy (anonymous):

-3 -6 = -9 - 3 ---------- = ---- = 3 6-9 = -3 -1

OpenStudy (anonymous):

why -3-6?

OpenStudy (anonymous):

its -3+-6

OpenStudy (anonymous):

Adding a negative is the same as subtracting.

OpenStudy (anonymous):

ooo wait i think i get it

OpenStudy (anonymous):

-9/6-9

OpenStudy (anonymous):

Yes 6-9 = -3. So -9 / -6 = 3

OpenStudy (anonymous):

o.

OpenStudy (anonymous):

-3* haha woops not -6

OpenStudy (anonymous):

ok but was i right in what i typed

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

-9/6-9?

OpenStudy (anonymous):

Yes, then simplify the denominator.

OpenStudy (anonymous):

um, do the answer =3?

OpenStudy (anonymous):

Yes